20170502 » History » Version 12
Nebrensky, Henry, 12 May 2017 18:49
minutes
1 | 1 | Boyd, Steven | h1. Ops/Comm Scheduling meeting : 2/5/2017 |
---|---|---|---|
2 | |||
3 | Location : CR21 at 14:00 pm |
||
4 | |||
5 | Phone access : *New* phonebridge number |
||
6 | UK: 08082380274 |
||
7 | US: 631.621.5253 |
||
8 | Access code: 1779339517 |
||
9 | |||
10 | h2. Agenda |
||
11 | |||
12 | * Actions |
||
13 | 9 | Boyd, Steven | ** -Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands- |
14 | 1 | Boyd, Steven | ** Macwaters - Re-make dodgy compressor2 remote cable connector - stands |
15 | 9 | Boyd, Steven | ** Macwaters - Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress |
16 | ** -Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress- |
||
17 | 1 | Boyd, Steven | ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress |
18 | ** Macwaters - Report on status of FC psu repair (stands) |
||
19 | *** *This needs to be cross-checked* - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it _in situ_ before the run? |
||
20 | 7 | Boyd, Steven | *** _Josef :"we just run 50A through it when it is connected to the magnet"_ |
21 | 1 | Boyd, Steven | ** Hall Manager - water system change-over procedure (stands) |
22 | ** EOverton - Status of @miceraid5@ disk cleanup (stands) |
||
23 | ** JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands) |
||
24 | 9 | Boyd, Steven | ** -CWhyte to contact DL to clarify their [[Hall_Task_List_Spring17|Hall Task List]] items- |
25 | 1 | Boyd, Steven | ** The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not |
26 | 7 | Boyd, Steven | *** _Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC_ |
27 | 9 | Boyd, Steven | ** -HNebrensky to send BLOC-chasing email- |
28 | 1 | Boyd, Steven | |
29 | * DC Update |
||
30 | * LH2 Update |
||
31 | 2 | Boyd, Steven | * Tracker Update |
32 | 1 | Boyd, Steven | * [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) |
33 | * Schedule for startup |
||
34 | 8 | Boyd, Steven | ** Startup run on Monday May 8th - Hall should be reserved for operations activities all day. Controlled Entry is possible |
35 | but only when permitted by the MOM. |
||
36 | 5 | Boyd, Steven | ** We will *not* be running the SS magnets this user cycle, or taking data with Neon. |
37 | ** We will need to run the FC at some point. |
||
38 | ** Co-ordination between DC and MOM will be key. |
||
39 | 1 | Boyd, Steven | |
40 | AOB: |
||
41 | |||
42 | * Last meeting : [[20170425|Tuesday Apr 25 2017]] |
||
43 | 10 | Nebrensky, Henry | * Next meeting : [[20170509|Tuesday May 9 2017]] |
44 | 11 | Nebrensky, Henry | |
45 | h3. Minutes |
||
46 | |||
47 | * Actions |
||
48 | ** -Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers - stands- |
||
49 | ** Macwaters - Re-make dodgy compressor2 remote cable connector - stands |
||
50 | ** Macwaters - Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress |
||
51 | ** -Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress- |
||
52 | ** Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress |
||
53 | ** Macwaters - Report on status of FC psu repair (stands) |
||
54 | *** *This needs to be cross-checked* - e.g. Josef Boehm to check with Trevor. What is plan when it does return - will we want to test it _in situ_ before the run? |
||
55 | *** _Josef :"we just run 50A through it when it is connected to the magnet"_ |
||
56 | *** Run into short in next couple of days |
||
57 | ** Hall Manager - water system change-over procedure (stands) |
||
58 | ** EOverton - Status of @miceraid5@ disk cleanup (stands) |
||
59 | ** JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands) |
||
60 | ** -CWhyte to contact DL to clarify their [[Hall_Task_List_Spring17|Hall Task List]] items- |
||
61 | ** The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not |
||
62 | *** _Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC_ |
||
63 | *** _Still need to decide, then get Trevor to re-connect if needed_ |
||
64 | ** -HNebrensky to send BLOC-chasing email- |
||
65 | |||
66 | * DC Update |
||
67 | Tracker cryo lost channels - possibly water ingress, looking for sample bottles. |
||
68 | DC to check that CC magnet supplies are all into shorts. |
||
69 | |||
70 | * LH2 Update |
||
71 | HAZOP on 16th/17th May; have the panel; hope to get LOPA done at end. |
||
72 | Working for the Neon test - will be "as LH2". Several days of cooling, and 1 day with liquid Neon. Quench test after run. |
||
73 | The LH2 Beacon will be live during the Neon test - should warn ISIS MCR beforehand. |
||
74 | |||
75 | * Tracker Update |
||
76 | One of the boards needs changing - Ed will do while he's down. |
||
77 | ** Tracker volume Helium purge: MU has added more putty round the leaky one. |
||
78 | When should we start the flow? Should estimate from volume volume and flowrate how long it takes to change the gas in the Tracker, then allow 5 volume changes before running. |
||
79 | Craig to show MOM how to start flow and change bottles |
||
80 | |||
81 | * [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) |
||
82 | |||
83 | * Schedule for startup |
||
84 | ** Startup run on Monday May 8th - Hall should be reserved for operations activities all day. Controlled Entry is possible |
||
85 | but only when permitted by the MOM. |
||
86 | ** We will *not* be running the CC magnets this user cycle, or taking data with Neon. |
||
87 | ** The PRY will stay *open*. |
||
88 | ** Co-ordination between DC and MOM will be key. |
||
89 | *** Run meetings 10 a.m. in ISIS coffee lounge |
||
90 | |||
91 | * AOB |
||
92 | 12 | Nebrensky, Henry | CRogers to check the [[Step4_Pionic_beam_Tags|M3 currents]] page and send to MOM. Also the CC current setting sign-off, even if it is null. |
93 | 1 | Boyd, Steven | CWhyte to request geometry survey of SS's and FC, and check TOF1 status. |
94 | 12 | Nebrensky, Henry | DS restart later this week |
95 | EO hopes to have @target1ctl@ finished by end Friday - get Joe down week after next. |
||
96 | Shifter training should always include some level of Controlled Entry refresher. |
||
97 | MOM is about to send off the Run Request to ISIS. |